# General Exploratory Bonus for Optimistic Exploration in RLHF

### v1

@@ D.1 Implementation Details
Following prior works Zhang et al. (2024a); Xie et al. (2024); Chen et al. (2025), we adopt the same algorithmic backbone for empirical validation to explicitly show the effect of different exploratory bonuses in the loss function. This algorithm bypasses the reward modeling in each iteration through reward reparameterization, known as iterative DPO (Dong et al., 2024). Previous works further reparameterize the bonus term to incorporate the algorithm. Since iterative DPO can seamlessly extend to f-divergence, we also follow prior works to reparameterize our general exploratory bonus. The detailed algorithm can be formulated as in Algorithm 1.
@@
All experiments are conducted on two NVIDIA H200 GPUs. When training and sampling, the max length is set to 2048. For training, the batch size per device is set to 2; we enable the gradient checkpointing and the gradient accumulation step is set to 64; the learning rate is 5e-7 with a cosine scheduler, and the warm-up ratio is 0.03. In the main experiments, we use the best performance with $\kappa$ with a suitable ratio range to f-dpo loss across 1,1e2,1e4,1e6,1e81,1e-2,1e-4,1e-6,1e-8. For sampling, the temperature is set to 1. For in-domain evaluation and MATH evaluation, we set temperature to 0.6 and top-p to 0.9; we use the default setting of alpaca-eval.
Appendix E Proofs of Lemma 4.1
Lemma 4.1
Denote two objectives as $h(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}uf^{\prime}(u)-f(u)$ and $\hat{h}(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{u}{Z(x)}f^{\prime}(u)-f(\frac{u}{Z(x)})$, where $u$ varies with $\pi$ and $\pi_{\textrm{ref}}$. When $f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z(x)})]/[Z(x)uf^{\prime\prime}(u)]=\Lambda(x)$ is constant in $y$ and $\Lambda(x)>0$, minimizing the two objectives $\min_{\pi}-h(u)$ and $\min_{\pi}-\hat{h}(u)$ induce the same class of policies.
Proof  For a fixed $x$, with the Lagrange multiplier and KKT conditions, when $\pi(y|x)>0$, we obtain
$\frac{\partial h}{\partial\pi(y|x)}$ $=u(x,y)f^{\prime\prime}(u(x,y))\cdot\frac{\partial\pi_{\textrm{ref}}(y|x)u(x,y)}{\partial\pi(y|x)}=\mu_{1}(x),$ (31)
where $\mu_{1}(x)$ is a dual variable with respective to $x$. When $\frac{f^{\prime}(u)+uf^{\prime\prime}(u)-f(\frac{u}{Z})}{Zuf^{\prime\prime}(u)}=\Lambda(x)$, we have
$\mu_{1}(x)\Lambda(x)=\frac{1}{Z}(f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z}))\cdot\frac{\partial\pi_{\textrm{ref}}(y|x)u(x,y)}{\partial\pi(y|x)},$ (32)
which directly follows by the KKT conditions of $\hat{h}(x)$, i.e. $\frac{\partial\hat{h}}{\partial\pi(y|x)}=\mu_{2}(x)$ where $\mu_{2}(x)$ is a dual variable equals to $\mu_{1}(x)\Lambda(x)$. Hence, every policy that satisfies the stationary condition for $h$ also satisfies it for $\hat{h}$. Since $\Lambda(x)>0$, the second-order derivative $\frac{\partial h}{\partial^{2}\pi(y|x)}$ and $\frac{\partial\hat{h}}{\partial^{2}\pi(y|x)}$ have the same sign, which indicates they share the same local minima. Hence, minimizing the two objectives $\min_{\pi}-h(u)$ and $\min_{\pi}-\hat{h}(u)$ induces the same class of policies.
$\square$

### v4

@@ C.1 Implementation Details
Following prior work (Zhang et al., 2024a; Xie et al., 2024; Chen et al., 2025), we adopt the same algorithmic backbone for empirical validation in order to isolate and compare the effects of different exploratory bonuses in the loss function. This backbone bypasses reward modeling at each iteration through reward reparameterization, a procedure commonly referred to as iterative DPO (Dong et al., 2024). Previous methods of exploratory bonuses further reparameterize their bonus terms to make them compatible with this framework. To extend the original framework to $\alpha$-divergence, we extend the iterative DPO with f-DPO (Wang et al., 2024) loss, and likewise reparameterize our generalized exploratory bonus accordingly. The full procedure is summarized in Algorithm 1. Since the only difference is the loss function in Line 5 of Algorithm 1, our GEB does not induce any additional compute costs.
@@
All experiments are conducted on two NVIDIA H200 GPUs. When training and sampling, the max length is set to 2048. For training, the batch size per device is set to 2; we enable gradient checkpointing, and the gradient accumulation step is set to 64; the learning rate is 5e-7 with a cosine scheduler, and the warm-up ratio is 0.03. In the main experiments, we use the best performance with $\kappa$ with a suitable ratio range to f-dpo loss across 1,1e2,1e4,1e6,1e81,1e-2,1e-4,1e-6,1e-8. For sampling, the temperature is set to 1. For in-domain evaluation and MATH evaluation, we set temperature to 0.6 and top-p to 0.9; we use the default setting of alpaca-eval.
Figure 4: Comparison on the bandit policy distributions trained with DPO (left) and GEB (right). The DPO policy collapses to a local optimum, while the GEB policy continues to explore and ultimately chooses the globally preferred action.
Figure 5: Initial reference bandit distribution (“ref”) and the reward distribution. Because the most preferred action lies in a low-probability region, it is rarely visited under purely passive exploration.
Appendix D Toy Experiments
To illustrate a setting in which GEB yields substantial improvement, we construct a toy example in which the most preferred action lies in a rarely visited region.
D.1 Experimental setting
We consider a 1000-arm bandit with 1000 parameters, each parameter corresponding to a distinct arm. As shown in Fig. 5, the most preferred action lies in a rarely visited region, making it unlikely to be sampled under pure passive exploration, as in f-DPO. Each experiment is run for 5000 iterations. At each iteration, the bandit policy generates 64 rollouts to form a batch of 32 preference pairs. The learning rate is set to 1e-2 with no warm-up phase.
D.2 Results and Analyses
As shown in Fig. 4, the bandit policy trained with DPO becomes trapped in a local optimum: its probability mass collapses onto a suboptimal action because the policy never encounters the truly preferred action during training. In contrast, all three GEB variants successfully recover the desired distribution (right panel of Fig. 4), concentrating probability on the most preferred action. This demonstrates that GEB effectively promotes exploration into low-probability regions, enabling the policy to discover and select the optimal action despite its small initial likelihood.
Appendix E Supplement results of main experiments
E.1 Repeated run of the main experiment
Table 7: Repeated in-domain evaluation on different exploration bonuses. Boldface and underline indicate the best and the second-best results, respectively. $\pi, GEB$-$1/\pi$, and GEB-$\mathrm{arctanh}(\pi-1)$ corresponds to $1+\alpha-\pi$, $1/\pi$, and $\mathrm{arctanh}(1-\pi)+\alpha$ as in Table 2.
KL ($\alpha$=1) Hel. ($\alpha$=0.5) f-KL ($\alpha$=0)
WR AvgR WR AvgR WR AvgR
Mistral-Instruct-v0.3
[…]
LLaMA-3-8B-SFT
[…]
Table 8: p-value of independent p-test between win-rates of GEB variants and f-DPO
f-DPO vs. Mistral-Instruct-v0.3 LlaMA-3-8B-SFT
KL. Hel. f-KL. KL. Hel. f-KL.
GEB-$\pi$ 0.003 0.014 0.034 0.145 0.084 0.117
GEB-$1/\pi$ 0.021 0.051 0.027 0.031 0.047 0.037
GEB-$\mathrm{arctanh}(\pi-1)$ 0.023 0.015 0.155 0.024 0.042 0.014
To assess the statistical significance of GEB’s performance gains, we repeat each experiment five times and report the mean and standard deviation. As shown in Table 7, GEB achieves consistently higher performance than baseline methods, with statistically significant improvements. Moreover, we conduct independent p-test between GEB variants and f-DPO. The p-value is shown in Table 8. Most p-values are less then 0.05, which means the improvement of GEB is statistically significant.
E.2 Semantic coherence of the sampled responses
We use GPT-4 to evaluate whether a given sentence is coherent, nonsensical, or contains meaningless content. The scoring scale ranges from 0 (fully coherent) to 3 (complete nonsense with no coherent meaning). We apply this evaluation to the responses generated in the final training iteration of DPO and the three GEB variants. The resulting scores are as follows.
Table 9: Semantic coherence scores of responses produced by policy models trained using DPO and the GEB variants.
DPO $\pi GEB$-$1/\pi$ GEB-$\text{arctanh}\pi$
KL. 1.24 1.08 1.32 1.19
Hel. 1.33 1.42 1.12 1.23
fKL. 1.32 1.38 1.32 1.30
As shown in Table 9, the responses produced by the GEB variants exhibit semantic coherence comparable to those generated by DPO. This indicates that, in practice, GEB promotes exploration into moderately underrepresented yet still semantically meaningful regions of the output space.
E.3 The choice of $u$
The three variants of the exploration bonuses in Table 2 represent different instantiations of the GEB framework. Each satisfies the optimism condition in Definition 3.1 and exhibits consistent performance improvements on the alignment task.
Nonetheless, the curvature of $u$ with respect to $\pi$ meaningfully affects the optimization dynamics of the exploratory bonus. For instance, when $u=1-\pi+\alpha$, the function is linear, thus the gradient of $u$ to $\pi$ is a constant. Therefore, the per-trajectory incentive to decrease $\pi$ is constant. In contrast, when $u$ is convex—such as $u=1/\pi$—the gradient magnitude diminishes as $\pi$ becomes larger. This results in a more conservative reduction of $\pi$ compared to the linear case.
Takeaway: The curvature of $u$ with respect to $\pi$ governs the behavior of the exploration bonus. Greater convexity leads to a more conservative shift of probability mass toward underexplored regions.
Table 10: Ablation study on the sample responses used for the bonus term calculation. We compare the results on UltraFeedback (win rate) across three divergences with all responses (all) and with the rejected responses only (rejected-only).
Method KL all KL rejected-only Hel. all Hel. rejected-only fKL all fKL rejected-only
GEB-$\pi$ 79.78 81.00 73.40 75.48 51.32 51.68
GEB-$\frac{1}{\pi}$ 79.56 80.00 72.25 73.97 51.61 52.26
GEB-$\mathrm{arctan}\pi$ 79.64 79.71 73.11 75.69 51.25 52.76
Restricting the exploratory bonus to rejected responses
Restricting the bonus term to rejected responses is a common practice in prior works on exploratory bonuses (Zhang et al., 2024a; Chen et al., 2025; Cen et al., 2025). Importantly, this restriction does not constitute a theoretical departure from our framework. The optimism guarantee in Theorem 4.2 hinges on increasing the probability of low $\pi_{\text{ref}}$ regions, i.e., underexplored regions. Because rejected samples lie precisely in these low-probability areas, applying the bonus only to rejected responses preserves the intended optimism direction. While preserving the theoretical guarantee on the optimism, it also shows a practical advantage as shown in Table 10.